Whirly Corporation’s contribution format income statement for the most recent month is shown below:

Total Per Unit
Sales (8,300 units) $ 282,200 $ 34.00
Variable expenses 157,700 19.00
Contribution margin 124,500 $ 15.00
Fixed expenses 54,600
Net operating income $ 69,900

Required:

(Consider each case independently):

1. What would be the revised net operating income per month if the sales volume increases by 40 units?

2. What would be the revised net operating income per month if the sales volume decreases by 40 units?

3. What would be the revised net operating income per month if the sales volume is 7,300 units?
Answer #1

ANSWER:

(1)

number of sales units = 8300 + 40 = 8340

revised net operating income = total contribution margin - fixed expenses

= (8340 x $15) - $54600

= $125100 - $54600

= $70500

(2)

number of sales units = 8300 - 40 = 8260

revised net operating income = total contribution margin - fixed expenses

= (8260 x $15) - $54600

= $123900 - $54600

= $69300

(3)

number of sales units = 7300

revised net operating income = total contribution margin - fixed expenses

= (7300 x $15) - $54600

= $109500 - $54600

= $54900